According to a report by Scarborough Research, the average monthly household cellular phone bill is $73. Suppose local monthly household cell phone bills are normally distributed with a standard deviation of $11.35.
a. What is the probability that a randomly selected monthly cell phone bill is more than $100?
b. What is the probability that a randomly selected monthly cell phone bill is between $60 and $83?
c. What is the probability that a randomly selected monthly cell phone bill is between $80 and $90?
d. What is the probability that a randomly selected monthly cell phone bill is no more than $55?
Show work

Answer by stanbon(75887)

According to a report by Scarborough Research, the average monthly household cellular phone bill is $73. Suppose local monthly household cell phone bills are normally distributed with a standard deviation of $11.35.
a. What is the probability that a randomly selected monthly cell phone bill is more than $100?
-------
z(100) = (100-73)/11.35 = 27/11.35 = 2.3789
P(x > 100) = P(z > 2.3789) = 0.0087
------------------------------------------
b. What is the probability that a randomly selected monthly cell phone bill is between $60 and $83?
z(83) = (83-73)/11.35 = 10/11.35 = 0.8811
z(60) = (60-73)/11.35 = -1.145
---
P(60< x <83) = P(-1.145< z < 0.8811) = 0.6848
-------------------------------------------
Use the same methods to work "C" and "D".
Cheers,
Stan H.
---------------
c. What is the probability that a randomly selected monthly cell phone bill is between $80 and $90?
d. What is the probability that a randomly selected monthly cell phone bill is no more than $55?